OpenStudy (anonymous):

Simplify completely, 4x^2 - 32x + 48 / 3x^2 - 17x - 6

OpenStudy (anonymous):

\[=\frac{ 4x^2-32x+48 }{ 3x^2-17x-6 }\] \[= \frac{ 4(x^2-8x+12) }{ 3x^2-17x-6 }\] \[= \frac{ 4(x-6)(x-2) }{ (x-6)(3x-1) }\] \[=\frac{ 4(x-2) }{ 3x-1 }\] there you go!

OpenStudy (anonymous):

THANKS YOU<3

OpenStudy (anonymous):

no problem ^^